OpenStudy (anonymous):

How would i find slope on a point on a graph? example (2,6)

OpenStudy (anonymous):

or (2,6) and (5,-3)

OpenStudy (anonymous):

That would be easy..

OpenStudy (anonymous):

\[\text{Slope = }\frac{\Delta y}{\Delta x}=\frac{ y_2-y_1}{ x_2-x_1}\]

OpenStudy (anonymous):

(-3)-6 over 5 - 3 =7.2? would be the slope ?

OpenStudy (anonymous):

No -3

OpenStudy (anonymous):

how

OpenStudy (anonymous):

-3-6/5-3=-9/3=-3

OpenStudy (anonymous):

whyd you divide by 3